### v3

@@ 5.1. Oriented exchange graphs are acyclic
If $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ is a green mutation, then $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subsetneq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. In particular, the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$ is acyclic.
@@
Since $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ is a green mutation and by Lemma 5.5, there exists
${\bf v}=(v_{1},\ldots,v_{n})^{T},\;{\bf v}^{\prime}=(v_{1}^{\prime},\ldots,v_{n}^{\prime})^{T}\in\mathbb{Z}_{>0}^{n}$
such that $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$. Set $u=\prod_{i=1}^{n}x_{i;t}^{v_{i}}$ and $u^{\prime}=\prod_{i=1}^{n}x_{i;t^{\prime}}^{v_{i}^{\prime}}$. Notice that we have
$({\bf g}_{u}^{t_{0}})^{\circ}=G_{t}{\bf v},\;\;({\bf g}_{u^{\prime}}^{t_{0}})^{\circ}=G_{t^{\prime}}{\bf v}^{\prime}.$
Since $\operatorname{\mathsf{supp}}(u)=[{\bf x}_{t}^{\circ}],\;\operatorname{\mathsf{supp}}(u^{\prime})=[{\bf x}_{t^{\prime}}^{\circ}]$ and by Corollary 5.4, we have that
$\operatorname{\mathsf{dom}}^{t_{0}}[t]$ $=$ $\operatorname{\mathsf{dom}}^{t_{0}}(u)=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{u}^{t_{0}}]=0\}=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[SG_{t}{\bf v}]=0\},$
$\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ $=$ $\operatorname{\mathsf{dom}}^{t_{0}}(u^{\prime})=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{t_{0}}]=0\}=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[SG_{t^{\prime}}{\bf v}^{\prime}]=0\}.$
@@
In rest part of the proof, we will show $x_{k;t^{\prime}}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ but $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}[t]$. Thus the inclusion $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ is strict.
By Proposition 5.3 (ii), we have $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$ and $[{\bf x}_{t^{\prime}}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. In particular,
@@
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}^{\prime}}^{t_{0}}]+F_{x_{k;t^{\prime}}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}]=F_{x_{k;t^{\prime}}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}].$
@@
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=0+F_{x_{k;t^{\prime}}}^{t}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t}]=(1+y_{k})[(S\mid{\bf 0}){\bf e}_{k}]=s_{k},$
where ${\bf e}_{k}$ is the $k$th column of $I_{m}$ and $s_{k}$ is the $(k,k)$-entry of $S=diag(s_{1},\ldots,s_{n})$. Thus
$F_{x_{k;t^{\prime}}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}]=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=s_{k}>0,$
@@
Take $\widetilde{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}=\Lambda$ and ${\bf x}=(x_{1},x_{2})$. Clearly, $\widetilde{B}^{T}\Lambda=(S\mid{\bf 0})=I_{2}$. The oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}=\mathcal{A}({\bf x},\widetilde{B})$ is as follows:
@@
We can see that the initial cluster variables $x_{1},x_{2}$ are contained in any dominant set with respect to the vertex $t_{0}$, as shown in Proposition 5.3 (i).
Let $\mathcal{H}_{0}$ be the set of seeds (up to seed equivalence) of $\mathcal{A}$, equivalently, $\mathcal{H}_{0}$ is the vertex set of the exchange graph of $\mathcal{A}$. For the rooted vertex $t_{0}\in\mathbb{T}_{n}$, we define an relation $\leq_{t_{0}}$ on $\mathcal{H}_{0}$ as follows: $[{\bf x}_{w},\widetilde{B}_{w}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$ if and only if there exists a sequence $\overleftarrow{\mu}$ of green mutations such that
$[\overleftarrow{\mu}({\bf x}_{w},\widetilde{B}_{w})]=[{\bf x}_{t},\widetilde{B}_{t}],$
i.e., there exists a path from the vertex $[{\bf x}_{w},\widetilde{B}_{w}]$ to the vertex $[{\bf x}_{t},\widetilde{B}_{t}]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$.
If $[{\bf x}_{w},\widetilde{B}_{w}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$ and $[{\bf x}_{t},\widetilde{B}_{t}]\neq[{\bf x}_{w},\widetilde{B}_{w}]$, we write $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$.
@@
The Hasse quiver of the poset set $(\mathcal{H}_{0},\leq_{t_{0}})$ is a quiver $\overrightarrow{\mathcal{H}}^{\prime}$ defined as follows:
  • the vertex set of $\mathcal{H}^{\prime}$ is $\mathcal{H}_{0}$;
  • $[{\bf x}_{w},\widetilde{B}_{w}]\rightarrow[{\bf x}_{t},\widetilde{B}_{t}]$ is an arrow in $\overrightarrow{\mathcal{H}}^{\prime}$ if and only if $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$ and there exists no seed $({\bf x}_{s},\widetilde{B}_{s})$ such that $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$.
Notice that the Hasse quiver $\overrightarrow{\mathcal{H}}^{\prime}$ and the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$ have the same vertex set $\mathcal{H}_{0}$. It is easy to see that the arrow set of $\overrightarrow{\mathcal{H}}^{\prime}$ is a subset of the arrow set of $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$. It is natural to ask when the two quivers coincide. The next subsection aims to answer this question.
@@ 5.2. Oriented exchange graphs are Hasse quivers
In this subsection, we will show that the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$ coincides with the Hasse quiver $\overrightarrow{\mathcal{H}}^{\prime}$ of the poset $(\mathcal{H}_{0},\leq_{t_{0}})$.
@@
Let $({\bf x}_{t},\widetilde{B}_{t})$ be a seed of $\mathcal{A}$. Denote by
$\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ $=$ $\{x\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\mid F_{z}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x}^{t_{0}}]=0,\;\forall z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\}$
@@
Then $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])=[{\bf x}_{t}^{\circ}]$ and in particular, $[{\bf x}_{t}]$ is uniquely determined by the dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$.
@@
So we have $x_{k;t}\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$. Since $k$ is an arbitrary integer $[1,n]$, we get $[{\bf x}_{t}^{\circ}]\subseteq\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$.
Now let us take $x\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$. We know that $F_{z}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x}^{t_{0}}]=0$ for any $z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]$. Since $x_{k;t}$ is in $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ for $k\in[1,n]$, we have $F_{x_{k;t}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x}^{t_{0}}]=0$ for $k\in[1,n]$.
Since $x\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$ and $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$, we have $x\in\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$, i.e., $F_{x}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}]=0$. Thus $(x_{k;t}\mid\mid x)_{F}=F_{x_{k;t}}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x}^{t_{0}}]+F_{x}^{t_{0}}[(S\mid{\bf 0}){\bf g}_{x_{k;t}}^{t_{0}}]=0$. Thus for any two cluster variables $z_{1},z_{2}$ in $[{\bf x}_{t}^{\circ}]\cup\{x\}$, we have $(z_{1}\mid\mid z_{2})_{F}=0$. Then by Lemma 4.21, we know that $[{\bf x}_{t}^{\circ}]\cup\{x\}$ is a subset of some cluster of $\mathcal{A}$. Since $[{\bf x}_{t}^{\circ}]$ is the set of unfrozen cluster variables in ${\bf x}_{t}$ and $x$ is an unfrozen cluster variable, we must have $x=x_{i;t}$ for some $i\in[1,n]$. Thus we get
$\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\subseteq[{\bf x}_{t}^{\circ}].$
Hence, $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])=[{\bf x}_{t}^{\circ}]$ and in particular, $[{\bf x}_{t}]$ is given by the union of $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ and the set of frozen variables. ∎
@@
Let $({\bf x}_{t},\widetilde{B}_{t})$, $({\bf x}_{s},\widetilde{B}_{s})$ and $({\bf x}_{w},\widetilde{B}_{w})$ be three seeds of $\mathcal{A}$. If
$[{\bf x}_{w},\widetilde{B}_{w}]\leq_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}],$
then $[{\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$.
@@
It suffices to prove that $[{\bf x}_{t}^{\circ}]\cap[{\bf x}_{w}^{\circ}]\subseteq[{\bf x}_{s}^{\circ}]$, because frozen variables are contained in all the clusters of $\mathcal{A}$.
Since $[{\bf x}_{w},\widetilde{B}_{w}]\leq_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]\leq_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$, we know that there is a sequence of green mutations from $[{\bf x}_{w},\widetilde{B}_{w}]$ to $[{\bf x}_{s},\widetilde{B}_{s}]$ and a sequence of green mutations from $[{\bf x}_{s},\widetilde{B}_{s}]$ to $[{\bf x}_{t},\widetilde{B}_{t}]$. Then by Theorem 5.7, we have $\operatorname{\mathsf{dom}}^{t_{0}}[w]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$.
By the definition of $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$, it is easy to see $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$. By Proposition 5.10, we know that $[{\bf x}_{t}^{\circ}]=\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ and $[{\bf x}_{s}^{\circ}]=\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$. Thus we have $[{\bf x}_{t}^{\circ}]\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq[{\bf x}_{s}^{\circ}]$.
By Proposition 5.3 (ii), we know that $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[w]$. Then by $\operatorname{\mathsf{dom}}^{t_{0}}[w]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]$, we have $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]$. So we know that $[{\bf x}_{t}^{\circ}]\cap[{\bf x}_{w}^{\circ}]\subseteq[{\bf x}_{t}^{\circ}]\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq[{\bf x}_{s}^{\circ}]$. This implies $[{\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$. ∎
@@
    There exists an arrow $[{\bf x}_{w},\widetilde{B}_{w}]\rightarrow[{\bf x}_{t},\widetilde{B}_{t}]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$.
@@
    $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}]$ and there exists no seed $({\bf x}_{s},\widetilde{B}_{s})$ such that
    $[{\bf x}_{w},\widetilde{B}_{w}]<_{t_{0}}[{\bf x}_{s},\widetilde{B}_{s}]<_{t_{0}}[{\bf x}_{t},\widetilde{B}_{t}].$

### v4

@@ 7.1. Oriented exchange graphs are acyclic
If $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ is a green mutation, then $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subsetneq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. In particular, the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$ is acyclic.
@@
Since $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ is a green mutation and by Lemma 7.6, there exist
${\bf v}=(v_{1},\ldots,v_{n})^{T},\;{\bf v}^{\prime}=(v_{1}^{\prime},\ldots,v_{n}^{\prime})^{T}\in\mathbb{Z}_{\geq 1}^{n}$
such that $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$. Set $u=\prod_{i=1}^{n}x_{i;t}^{v_{i}}$ and $u^{\prime}=\prod_{i=1}^{n}x_{i;t^{\prime}}^{v_{i}^{\prime}}$. We know that
${\bf g}_{u}^{t_{0}}=G_{t}{\bf v},\;\;{\bf g}_{u^{\prime}}^{t_{0}}=G_{t^{\prime}}{\bf v}^{\prime}.$
Since $\operatorname{\mathsf{supp}}(u)=[{\bf x}_{t}],\;\operatorname{\mathsf{supp}}(u^{\prime})=[{\bf x}_{t^{\prime}}]$ and by Corollary 7.5, we have that
$\operatorname{\mathsf{dom}}^{t_{0}}[t]$ $=\operatorname{\mathsf{dom}}^{t_{0}}(u)=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[S{\bf g}_{u}^{t_{0}}]=0\}=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[SG_{t}{\bf v}]=0\},$
$\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ $=\operatorname{\mathsf{dom}}^{t_{0}}(u^{\prime})=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[S{\bf g}_{u^{\prime}}^{t_{0}}]=0\}=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[SG_{t^{\prime}}{\bf v}^{\prime}]=0\}.$
@@
In the rest part of the proof, we will show $x_{k;t^{\prime}}\in\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ but $x_{k;t^{\prime}}\notin\operatorname{\mathsf{dom}}^{t_{0}}[t]$. Thus the inclusion $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$ is strict.
By Proposition 7.4 (ii), we have ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$ and ${\bf x}_{t^{\prime}}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t^{\prime}]$. In particular,
@@
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t_{0}}[S{\bf g}_{x_{k;t}^{\prime}}^{t_{0}}]+F_{x_{k;t^{\prime}}}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}]=F_{x_{k;t^{\prime}}}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}].$
@@
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=0+F_{x_{k;t^{\prime}}}^{t}[S{\bf g}_{x_{k;t}}^{t}]=(1+y_{k})[S{\bf e}_{k}]=s_{k},$
where ${\bf e}_{k}$ is the $k$th column of $I_{n}$ and $s_{k}$ is the $(k,k)$-entry of $S=diag(s_{1},\ldots,s_{n})$. Thus
$F_{x_{k;t^{\prime}}}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}]=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=s_{k}>0,$
@@
Take $B=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$, ${\bf x}=(x_{1},x_{2})$ and $S=I_{2}$. The oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}=\mathcal{A}({\bf x},B)$ is as follows:
@@
We can see that the initial cluster variables $x_{1},x_{2}$ are contained in any dominant set with respect to the vertex $t_{0}$, this is because the $F$-polynomials of initial cluster variables are $1$.
Let $\mathcal{H}_{0}$ be the set of seeds (up to seed equivalence) of $\mathcal{A}$, equivalently, $\mathcal{H}_{0}$ is the vertex set of the exchange graph of $\mathcal{A}$. For the rooted vertex $t_{0}\in\mathbb{T}_{n}$, we define an relation $\leq_{t_{0}}$ on $\mathcal{H}_{0}$ as follows: ${\bf x}_{w},B_{w}]\leq_{t_{0}}[{\bf x}_{t},B_{t}]$ if and only if there exists a sequence $\overleftarrow{\mu}$ of green mutations such that
$\overleftarrow{\mu}({\bf x}_{w},B_{w})]=[{\bf x}_{t},B_{t}],$
i.e., there exists a path from the vertex ${\bf x}_{w},B_{w}]$ to the vertex ${\bf x}_{t},B_{t}]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ of $\mathcal{A}$.
If ${\bf x}_{w},B_{w}]\leq_{t_{0}}[{\bf x}_{t},B_{t}]$ and ${\bf x}_{t},B_{t}]\neq[{\bf x}_{w},B_{w}]$, we write ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{t},B_{t}]$.
@@
[…]
@@ 7.2. Oriented exchange graphs are Hasse quivers
[…]
@@
Let $({\bf x}_{t},B_{t})$ be a seed of $\mathcal{A}$. Denote by
$\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ $=\{x\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\mid F_{z}^{t_{0}}[S{\bf g}_{x}^{t_{0}}]=0,\;\forall z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]\}$
@@
Then $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])=[{\bf x}_{t}]$ and in particular, ${\bf x}_{t}]$ is uniquely determined by the dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$.
@@
So we have $x_{k;t}\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$. Since $k$ is an arbitrary integer $[1,n]$, we get ${\bf x}_{t}]\subseteq\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$.
Now let us take $x\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$. We know that $F_{z}^{t_{0}}[S{\bf g}_{x}^{t_{0}}]=0$ for any $z\in\operatorname{\mathsf{dom}}^{t_{0}}[t]$. Since $x_{k;t}$ is in $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ for $k\in[1,n]$, we have $F_{x_{k;t}}^{t_{0}}[S{\bf g}_{x}^{t_{0}}]=0$ for $k\in[1,n]$.
Since $x\in\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}(x_{k;t})$, we have $F_{x}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}]=0$. Thus
$(x_{k;t}\mid\mid x)_{F}=F_{x_{k;t}}^{t_{0}}[S{\bf g}_{x}^{t_{0}}]+F_{x}^{t_{0}}[S{\bf g}_{x_{k;t}}^{t_{0}}]=0.$
Thus for any two cluster variables $z_{1},z_{2}$ in ${\bf x}_{t}]\cup\{x\}$, we have $(z_{1}\mid\mid z_{2})_{F}=0$. Then by Lemma 4.18, we know that ${\bf x}_{t}]\cup\{x\}$ is a subset of some cluster of $\mathcal{A}$. We must have $x=x_{i;t}$ for some $i\in[1,n]$. Thus we get
$\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\subseteq[{\bf x}_{t}].$
Hence, $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])=[{\bf x}_{t}]$. ∎
@@
Let $({\bf x}_{t},B_{t})$, $({\bf x}_{s},B_{s})$ and $({\bf x}_{w},B_{w})$ be three seeds of $\mathcal{A}$. If
$\operatorname{\mathsf{dom}}^{t_{0}}[w]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t],$
then ${\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{s}]$.
@@
By the definition of $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$, we can see $\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$. By Proposition 7.11, we know that ${\bf x}_{t}]=\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[t])$ and ${\bf x}_{s}]=\mathcal{P}(\operatorname{\mathsf{dom}}^{t_{0}}[s])$. Thus we have ${\bf x}_{t}]\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq[{\bf x}_{s}]$.
By Proposition 7.4 (ii), we know that ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[w]$. Then by $\operatorname{\mathsf{dom}}^{t_{0}}[w]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]$, we have ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]$. So we know that ${\bf x}_{t}]\cap[{\bf x}_{w}]\subseteq[{\bf x}_{t}]\cap\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq[{\bf x}_{s}]$. ∎
@@
    There exists an arrow ${\bf x}_{w},B_{w}]\rightarrow[{\bf x}_{t},B_{t}]$ in the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$.
@@
    ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{t},B_{t}]$ and there exists no seed $({\bf x}_{s},B_{s})$ such that
    ${\bf x}_{w},B_{w}]<_{t_{0}}[{\bf x}_{s},B_{s}]<_{t_{0}}[{\bf x}_{t},B_{t}].$